### v1

@@ 3.1 Hypercomplex Dense layer
We start with a description of the dense layer. It is a general-purpose layer that operates on the data with additional dimensionality, which is a multiple algebra dimensions. We assume that the input data are of dimension $b\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $al$ the algebra size, and $in$ the positive integer multiplier. The last two numbers determine the input data size. The input tensor $X=X_{i_{b}Rs(i_{al},i_{in})}$, where $i_{b}$ batch index, $i_{al}$ is the dimension of algebra, $i_{in}$ is the multiplicity index of algebra dimension. Moreover, we use learning parameters (weights/kernel) $K=K_{i_{al}i_{in}i_{u}}$, where $i_{u}$ is the index over units/neurons. The bias $b=b_{Rs(i_{al},i_{u})}$ is used if needed. Kernel and bias are usually initialized with numbers taken from specific distributions [9].
We now provide the algorithm of hypercomplex dense network in Algorithm 1. We offer both tensorial and abstract index notations (AIN). We need two flags bias - if bias is included and activation - if activation function $\sigma$ is used.
@@
$Output\leftarrow C_{1,0}(X,W)$ [AIN: $Outputs_{i_{b}i_{2}}\leftarrow\sum_{k}X_{i_{b}k=Rs(i_{al},i_{in})}W_{ki_{2}}$]
@@
The Keras with TensorFlow implementation and PyTorch implementations are given in [13].
@@ 3.2 Hypercomplex Convolutional layer
The additional ’image channels’ of the data can be packed as an element of algebra. For instance, the two-dimensional image can be decomposed as a matrix of four-dimensional algebra elements, i.e., color channels plus alpha, making a single pixel an element of four-dimensional algebra.
The general idea of NN action, as in [18], is to use algebra constants $A$ tensor to separate the coefficients of the algebra base component and then apply the traditional convolution for each algebra component.
The dimension of the input data $X=X_{i_{b}i_{1}\ldots i_{k}Rs(i_{al},i_{in})}$ of dimension $b\times n_{1}\times\ldots\times n_{k}\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $n_{i}$, $i\in\{1,\ldots,k\}$ the size of data sample in each dimension, $al$ the algebra dimension, and $in$ the positive multiplier [^3] One can see that the size of the data sample in each dimension (axes) stands right after the batch size, which is used in TensorFlow convention. For PyTorch, this multiindex is placed at the end. We will not provide an implementation for the PyTorch convention since it differs by a few permutations.. The kernel size is $al\times L_{1}\times\ldots L_{k}\times in\times F$, where: $L_{i}$, $i\in\{1,\ldots,k\}$ is the kernel dimension in each dimension, and $F$ the number of filters. We therefore have the kernel $K=K_{i_{al}i_{l1}\ldots i_{lk}i_{in}i_{f}}$. The flag $bias$ indicates the bias $b$ usage. If it is used, it has dimension $al\times F$. Kernel and bias can be initialized by arbitrary distributions [9]. We use standard $k$-dimensional convolution $convkD(X,K,strides,padding)$ for convoluting algebra components, there are standard optimized convolution operations [12]. The algorithm is presented in the Algorithm 2.
@@
$W\leftarrow C_{1,0}(A,K)$ [AIN: $W_{i_{a}k_{a}i_{l1}\ldots l_{lk}i_{in}i_{f}}\leftarrow\sum_{j}A_{i_{a}jk_{a}}K_{ji_{l1}\ldots i_{lk}i_{in}i_{f}}$]
$W\leftarrow\tau_{p=(0\rightarrow k+2)}(W)$ [AIN: $W_{k_{a}i_{l1}\ldots l_{lk}i_{a}i_{in}i_{f}}\leftarrow W_{p(i_{a}k_{a}i_{l1}\ldots l_{lk}i_{in}i_{f})}$]
$W\leftarrow Rs_{k+2}(W)$ [AIN: $W_{k_{a}i_{l1}\ldots l_{lk}ji_{f}}\leftarrow W_{k_{a}i_{l1}\ldots l_{lk}\underbrace{Rs(i_{a},i_{in})}_{=j}i_{f}}$]
@@ 4 Conclusion
In this paper, we introduced the mathematical details of the implementation of dense and convolutional NN based on hypercomplex and general algebras. The critical point in this presentation is to associate algebra multiplication with rank-three tensor. Thanks to this observation, all the NN processing steps can be represented as tensorial operations.
The fully tensorial operations applied to algebra operations simplifies neural networks operations, and allows to support for fast tensorial operations in modern packages as TensorFlow and PyTorch. Implementation of the above generalized hypercomplex NN is described elsewhere [13].
@@
[…]
Acknowledgements This paper has been supported by the Polish National Agency for Academic Exchange Strategic Partnership Programme under Grant No. BPI/PST/2021/1/00031 (nawa.gov.pl).

### v2

@@ 3.1 Hypercomplex dense layer
We start with a description of the dense hypercomplex layer. It is a general-purpose layer that operates on the data with additional dimensionality/axis, which is a multiple of algebra dimensions. We assume that the input data are of dimension $b\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $al$ the algebra size, and $in$ the positive integer multiplier. The last two numbers determine the input data size $n$. The input tensor in abstract index notation is $X=X_{i_{b}Rs(i_{al},i_{in})}$, where $i_{b}$ batch index, $i_{al}$ is the algebra index, $i_{in}$ is the multiplicity index of algebra dimension. Moreover, we use learning parameters (weights/kernel) $K=K_{i_{al}i_{in}i_{u}}$, where $i_{u}$ is the index over units/neurons. The bias $b=b_{Rs(i_{al},i_{u})}$ is used if needed. Kernel and bias are usually initialized with numbers taken from specific distributions; cf. [10].
The algorithm of the hypercomplex dense layer is provided in Algorithm 1. We offer both tensorial and abstract index notations (AIN). Since indices are in the subscript, so we write the summation sign explicitly. We need two flags bias - if bias is included and activation - if activation function $\sigma$ is used.
@@
$Output\leftarrow C_{1,0}(X,W)$ [AIN: $Output_{i_{b}i_{2}}\leftarrow\sum_{k}X_{i_{b}k=Rs(i_{al},i_{in})}W_{ki_{2}}$]
@@
The implementation using Keras with TensorFlow library implementation and PyTorch library implementations are described in [15] and implemented in Hypercomplex Keras library [22] in the HyperDense class. As it was emphasized, for $4$-diemnsional algebras the output agrees with the results of [20].
@@ 3.2 Hypercomplex convolutional layer
The additional ’image channels’ of the data (typically $3$-dimensional subarray of RGB colors) can be packed as an element of algebra. For instance, the two-dimensional image can be decomposed as a matrix of four-dimensional algebra elements, i.e., color channels plus alpha value, making a single pixel an element of four-dimensional algebra.
The general idea of the NN operation is to use algebra multiplication matrix tensor $A$ to separate the coefficients of the algebra base component and then apply the traditional convolution for each algebra component. This idea is an extension of the results of [20] for four-dimensional algebras, and in particular, of [9] for Quaternions. In [20], the split of components of algebra was implemented at the level of a specially tailored multiplication matrix, so it was not general.
The dimension of the input data $X=X_{i_{b}i_{1}\ldots i_{k}Rs(i_{al},i_{in})}$ of dimension $b\times n_{1}\times\ldots\times n_{k}\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $n_{i}$, $i\in\{1,\ldots,k\}$ the size of data sample in each dimension (e.g., for $k=2$ the raster coordinates $i_{1}$, $i_{2}$), $al$ the algebra dimension, and $in$ the positive multiplier [^3] One can see that the size of the data sample in each dimension (axes) stands right after the batch size, which is used in TensorFlow convention. For PyTorch, this multiindex is placed at the end. We will not provide an implementation for the PyTorch convention since it differs by a few permutations.. The kernel size is $al\times L_{1}\times\ldots L_{k}\times in\times F$, where: $L_{i}$, $i\in\{1,\ldots,k\}$ is the kernel dimension for each axis, and $F$ the number of filters. We therefore have the kernel $K=K_{i_{al}i_{l_{1}}\ldots i_{l_{k}}i_{in}i_{f}}$. The flag $bias$ indicates the bias $b$ usage. If it is used, it has dimension $al\times F$. Kernel and bias can be initialized by arbitrary distributions; see [10]. We use standard $k$-dimensional convolution $convkD(X,K,strides,padding)$ for convoluting algebra components, there are standard optimized convolution operations [14]. The algorithm is presented in the Algorithm 2.
@@
$W\leftarrow C_{1,0}(A,K)$ [AIN: $W_{i_{a}k_{a}i_{l_{1}}\ldots l_{l_{k}}i_{in}i_{f}}\leftarrow\sum_{j}A_{i_{a}jk_{a}}K_{ji_{l_{1}}\ldots i_{l_{k}}i_{in}i_{f}}$]
$W\leftarrow\tau_{p=(0\rightarrow k+2)}(W)$ [AIN: $W_{k_{a}i_{l1}\ldots l_{lk}i_{a}i_{in}i_{f}}\leftarrow W_{p(i_{a}k_{a}i_{l_{1}}\ldots l_{l_{k}}i_{in}i_{f})}$]
$W\leftarrow Rs_{k+2}(W)$ [AIN: $W_{k_{a}i_{l_{1}}\ldots l_{l_{k}}ji_{f}}\leftarrow W_{k_{a}i_{l_{1}}\ldots l_{l_{k}}\underbrace{Rs(i_{a},i_{in})}_{=j}i_{f}}$]
@@
The algorithm for different $k\in\{1,2,3\}$ is implemented in the classes HyperConv1D, HyperConv2D, and HyperConv3D of Hypercomplex Keras library [22] and described in [15]. It agrees and extends the results of [20] from $4$-dimensional algebra to arbitrary dimensional algebras.
@@ 4 Conclusion
We introduced the mathematical details of the implementation of dense and convolutional NN based on hypercomplex and general algebras. The critical point in this presentation is to associate algebra multiplication with rank-three tensor. Thanks to this observation, all the NN processing steps can be represented as tensorial operations.
The fully tensorial operations applied to algebra operations simplify neural network operations and allow support for fast tensorial operations in modern packages such as TensorFlow and PyTorch. Implementation of the above generalized hypercomplex NN is described elsewhere [15].
It is an extension of algorithms presented for $4$-dimensional algebras in [20] and for Quaternions in [9]. They are implemented in Hypercomplex Keras library [22] and checked to agree for $4$-dimensional algebras with the results from [20].